Write an equation in point-slope form for the line through the given points or through the given point with the given slope. (-2,3); m = -1
Answer by butterfliiizzz(18)

y – y1 = m (x – x1)
Whenever you have a point, it is in the form: (x, y)
When inserting a point into the point slope equation, you will put it where the x1 and the y1 are located, because those subscripts are indicators of a SPECIFIC point.
So, given the point (-2, 3) and the slope m = -1, our equation will look like this:
y – 3 = -1 (x - -2)

When you have minus a negative number, it becomes a positive or a plus:
y – 3 = -1 (x + 2)
Now we must distribute the -1 by multiplication:
y – 3 = -x – 2
We want this in y = mx + b format, so we must get the y by itself by adding 3 to each side:
y = -x +1 is the answer.